# Dual Alignment Maximin Optimization for Offline Model-based RL

### v2

@@ 6.2 Comparison Results
Dataset. We evaluate DAMO’s performance using the D4RL benchmark (Fu et al., 2020) within the MuJoCo simulation environment. Our comprehensive assessment covers 12 distinct datasets across three environments (HalfCheetah, Hopper, and Walker2d) and four dataset categories: random, medium, medium-replay, and medium-expert. Following standard evaluation protocols, we exclusively utilize the ”v2” versions of all datasets for consistent comparison.
@@
Baseline. We compare DAMO with two types of offline RL baselines: DICE-based algorithms which follow a similar form of training objective with DAMO and Model-based algorithms which achieve SOTA performance. For DICE-based approaches, Algae-DICE (Nachum et al., 2019b) transforms the intractable state-action-level constraint into a unified objective for policy training. Opti-DICE (Lee et al., 2021) directly estimates the stationary distribution corrections of the optimal policy to attain a high-rewarding policy. O-DICE (Mao et al., 2024a) using the orthogonal-gradient update to diminish the confliction of different gradients during policy training. For model-based approaches, MOPO (Yu et al., 2020) penalizes rewards via the predicted variance of ensembled dynamics models. COMBO (Yu et al., 2021) implements CQL within a model-based framework. TT (Janner et al., 2019) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter et al., 2022) adversarially trains the policy and the dynamics model within a robust framework. MOBILE (Sun et al., 2023) penalizes the value learning of synthetic data based on the estimated uncertainty of the Bellman Q-function, derived from an ensemble of models.
@@
Comparison Results. Table 2 reports the scores in D4RL benchmark. Overall, DAMO demonstrates superior performance, achieving the highest average score across all baseline methods. Notably, for high-quality datasets such as medium-expert and medium-replay, DAMO consistently attains state-of-the-art (SOTA) or near-SOTA performance. These results suggest that effective OOD constraints, combined with a high-quality behavior policy, enable the agent to learn high-reward policies within in-distribution (ID) regions. Furthermore, DAMO maintains strong performance on medium-quality datasets. This demonstrates the robustness of DAMO in handling less-than-ideal data conditions and highlights the importance of ensuring both model-environment policy consistency and synthetic-offline data compatibility. However, the performance of DAMO on random datasets reveals limitations, indicating the need for more aggressive exploration strategies in low-quality data scenarios to identify regions beyond those covered by the offline dataset. This phenomenon is particularly pronounced in the Walker2d-random dataset, where DAMO struggles to learn an effective policy. One possible explanation is that the optimization landscape of the environment may contain numerous suboptimal solutions corresponding to saddle points in the maximin framework. If this is the case, such structural characteristics could potentially make DAMO more susceptible to becoming trapped in local optima, thereby limiting its ability to discover better solutions. However, further investigation is needed to confirm this hypothesis.
6.3 Effects of Coefficient $\alpha$
@@
Figure 5: Two sets of experiments on hyperparameter $\alpha$. (a) The tuning experiment of $\alpha$ was performed on the hopper-medium-expert dataset. (b) The fixing experiment of $\alpha$ was performed on the halfcheetah-medium-expert dataset for validation.
@@
DAMO is regulated by a key hyperparameter: the hyperparameter $\alpha$ which controls the extent of policy discrepancies. In practical implementation, we fix $\alpha$ in the actor training. While we consider $\alpha$ as a hyperparameter controlling the degree of conservatism, the empirical results in Fig. 5(a) demonstrate that DAMO maintains robust performance across a wide range of $\alpha$ values, with its primary impact observed in the final stage performance. Notably, DAMO achieves superior performance with larger $\alpha$ values (e.g., $\alpha=3.0,5.0$), whereas insufficiently small values (e.g., $\alpha=1.0$) tend to promote excessive risk-taking during exploration, ultimately failing to converge to an optimal policy. To enhance training stability, we initially fix $\alpha$ to 1.0 in the actor training. However, experimental results demonstrate that DAMO achieves competitive performance even without this specific implementation. To further validate this observation, we conduct comparative experiments on the HalfCheetah environment, evaluating both the standard DAMO implementation and its variant without fixed $\alpha$ under their respective optimal hyperparameter configurations. As shown in Fig. 5(b), both settings achieve a competitive performance, confirming that DAMO maintains robust performance regardless of the implementation of fixing $\alpha$.
7 Conclusion and Discussion
@@
This paper critically examines limitations in existing offline model-based RL approaches, highlighting the crucial need to address behavior inconsistency caused by policy discrepancies between biased models and real environments. Building upon our analysis, we introduce a unified framework that concurrently ensures model-environment policy consistency and synthetic-offline data compatibility, with comprehensive experimental validation demonstrating its effectiveness. Significantly, the maximin optimization framework of DAMO demonstrates the benefits of aligning the actor and critic training objectives within the actor-critic architecture, ensuring essential conservatism during policy updates. The limitations of our work are focused on two-fold: the sampling of the initial state and the computing of the data alignment term. Although we assume the distribution of the initial state is known, the offline dataset does not explicitly label which states are initial states. Therefore, we randomly sample a state from the offline dataset to serve as the initial state, which creates a gap between theory and practice. When calculating the data alignment term, we trained a classifier to assist in the process. However, the classifier’s accuracy is critical: imprecise classifiers fail to estimate the modified reward accurately, leading to unstable training and potentially invalidating the practical objective as a lower bound. Future work could extend DAMO to areas such as offline-to-online RL (Li et al., 2023), where methods with strong adaptability and the ability to bridge the synthetic-to-real distribution gap are in high demand.
Impact Statement
This work presents an effective solution for addressing the distribution shift in offline model-based RL through dual policy-environment alignment, enabling robust simulation-to-reality transfer in safety-critical domains such as medical robotics. While advancing certifiable policy deployment via implicit behavior constraints and adaptive value estimation, challenges remain in reward specification sensitivity and residual model bias mitigation. The framework highlights the critical need for continuous dynamic alignment calibration to ensure reliable real-world RL system performance.
@@ Appendix A Proofs in the Main Text
We present a detailed proof process for each theorem discussed in the paper. The sequence of proofs follows the logic of derivation rather than strictly adhering to their order of appearance in the main text.
Theorem A.1.
The KL-divergence has the following upper bound:
$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})\leq\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
@@
$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ $=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]$
$=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}\cdot\dfrac{\rho_{M}^{\pi}}{\rho_{M}^{\pi}}]$
$=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
$\leq\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
Corollary A.2.
We have the following surrogate objective
$\arg\max\limits_{\pi}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
A discounted factor $\alpha$ will be added for training stability.
$\arg\max\limits_{\pi}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ (5)
Lemma A.3.
We can represent f-divergence in a function form: $D_{f}(x\|p)=\mathbb{E}_{z\sim p}f(\frac{x}{p})$, then its Fenchel conjugate function is $\mathbb{E}_{z\sim p}f_{\star}(y(z))$
@@
We use $g_{\star}(x)$ to denote $D_{f}(x\|p)$
$g_{\star}(x)$ $=\sup\limits_{y\in\mathcal{X}}\{\langle x,y\rangle-D_{f}(y\|p)\}$
[…]
$=\int_{R}p(z)\sup\limits_{z}[x(z)\frac{y(z)}{p(z)}-f(\frac{y(z)}{p(z)})]dz$
$=\int_{R}p(z)\sup\limits_{z}[x(z)y(z)-f(y(z))]dz$
[…]
$=\mathrm{E}_{z\sim p}f_{\star}(y(z))$
Corollary A.4.
For f-divergence $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})=\max\limits_{y(s,a,s^{\prime})}\mathrm{E}_{\rho_{T}^{\pi}}[y(s,a,s^{\prime})]-\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
@@
Verifying $D_{f}$ is a convex, closed function. A convex and closed function’s double conjugate function is itself, so $D_{f}$ can be represented in a MAX form. ∎
Theorem A.5.
Surrogate objective (5) is equivalent to the following maximin problem
$\max\limits_{\pi}\min\limits_{y(s,a,s^{\prime})}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]+\alpha\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
@@
According to lemma A.4, the objective can be presented in a Min form. ∎
Theorem A.6.
The following objective is equivalent to the surrogate objective
$\max\limits_{\pi}\min\limits_{Q(s,a)}J(\pi,Q(s,a))=\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\mathrm{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$ (6)
@@
Defining $y(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)}{\alpha}$. For the convenience, we define $\beta_{t}(s,a,s^{\prime})$
@@
Then, $\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}$. Similarly, we can define $\beta_{t}(s,a,s^{\prime},a^{\prime})$.
Substituting $y(s,a,s^{\prime})$ with $[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)]/\alpha$ and we have:
$\max\limits_{\pi}\min\limits_{Q(s,a)}$ $\mathrm{E}_{\rho_{T}^{\pi}}[Q(s,a)-\tau^{\pi}Q(s,a)]$
@@
$=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q(s,a)-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
$=$ $(1-\gamma)[\sum\limits_{s,a}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a)Q(s,a)-\sum\limits_{s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$

### v3

@@ 6.2.1. D4RL
Dataset. We first evaluate the performance of DAMO using the D4RL Fu et al. (2020). Our assessments cover 12 distinct datasets across three control environments (HalfCheetah, Hopper, and Walker2d) and four dataset categories: random, medium, medium-replay, and medium-expert. Following standard evaluation protocols, we utilize the ”v2” versions of all datasets in D4RL for consistent comparison.
@@
Baseline. We compare DAMO with two types of offline RL baselines: DICE-based algorithms, which follow a similar form of training objective to DAMO, and Model-based algorithms, which achieve SOTA performance in D4RL benchmark. For DICE-based approaches, Algae-DICE Nachum et al. (2019b) transforms the intractable state-action-level constraint into a unified objective for policy training. Opti-DICE Lee et al. (2021) directly estimates the stationary distribution corrections of the optimal policy to attain a high-rewarding policy. O-DICE Mao et al. (2024b) uses the orthogonal gradient update to diminish the conflict of different gradients during policy training. For model-based approaches, MOPO Yu et al. (2020) penalizes rewards via the predicted variance of ensembled dynamics models. COMBO Yu et al. (2021) implements CQL within a model-based framework. TT Janner et al. (2019) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO Rigter et al. (2022) adversarially trains the policy and the dynamics model within a robust framework. MOBILE Sun et al. (2023) penalizes the value learning of synthetic data based on the estimated uncertainty of the Bellman Q-function, which is derived from an ensemble of models.
@@
Comparison Results. Table 1 reports the scores in the D4RL benchmark. Overall, DAMO demonstrates superior performance, achieving the highest average score across all baseline methods. Notably, for high-quality datasets such as medium-expert and medium-replay, DAMO consistently attains SOTA or near-SOTA performance. These results suggest that effective dual alignment, combined with a high-quality behavior policy, enable the agent to learn high-reward policies within ID regions. Furthermore, DAMO maintains strong performance on medium-quality datasets. This demonstrates the robustness of DAMO in handling less-than-ideal data conditions and highlights the importance of ensuring both model-environment policy consistency and synthetic-offline data compatibility. However, the performance of DAMO on random datasets reveals limitations to some extent, indicating the need for more aggressive exploration strategies in low-quality data scenarios to identify uncovered regions. This phenomenon is pronounced in the Walker2d-random dataset, where DAMO struggles to learn an effective policy. One possible explanation is that the optimization landscape of the Walker2d contains numerous suboptimal solutions corresponding to saddle points in the maximin problem. If this is the case, such structural characteristics could potentially trap DAMO in local optima, limiting its ability to discover better solutions. Further investigation is needed to confirm the above hypothesis.
6.2.2. NeoRL
Table 4. Average returns on halfcheetah-jump. Batch Mean and Batch Max denote the mean and maximum average return across all trajectories in the offline dataset, respectively.
Environment Batch Mean Batch Max DAMO MOPO COMBO BEAR BRAC-v TD3+BC
halfcheetah-jump -1184 2887 $4871\pm 416$ $5030\pm 435$ $4868\pm 175$ $186\pm 231$ $1038\pm 179$ $1340\pm 303$
Dataset. To further assess the performance of DAMO in real-world scenarios, we test it in the NeoRL Qin et al. (2022). By emulating real-world data-collection scenarios, the datasets in NeoRL characterize a narrow scope and limited coverage, which impose significant challenges for offline RL. Our assessments include nine datasets across three environments (HalfCheetah, Hopper, Walker2d) and three dataset quality levels (L, M, H), corresponding to low, medium, and high.
Baseline. We have not chosen the DICE-based method for comparison, since the results of these methods can not be found in the original paper or the NeoRL paper, and searching for optimal hyperparameters for each method would take excessive time. Instead, we select SOTA offline model-free methods to ensure the diversity of the baseline. For model-based approaches, we keep two representative methods, MOPO Yu et al. (2020) and MOBILE Sun et al. (2023) for comparison. For model-free approaches, BC Bain and Sammut (1995) imitates the behavior policy from the offline dataset. CQL Kumar et al. (2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC Fujimoto and Gu (2021) introduces a behavior cloning regularization into the TD3 objective. EDAC An et al. (2021) penalizes the Q-function based on uncertainty derived from an ensemble of Q-networks.
Comparison Results. Table 3 reports the scores in the NeoRL benchmark. DAMO maintains competitive performance under more realistic and difficult scenarios, indicating strong potential for real-world applicability. Specifically, the strong generalization ability in narrow data coverage induced by suboptimal policy facilitates DAMO in achieving SOTA or near-SOTA performance on datasets of low-quality and medium-quality. Among high-quality datasets, DAMO secures a near-SOTA score in the Walker2d-H dataset and a high ranking score in the Hopper-H dataset. However, while training with the Halfcheetah-H dataset, DAMO fails to learn the optimal in-distribution policy, which contradicts our analysis in D4RL results. Through inspecting the training process, we observe that although DAMO achieves rapid policy improvement during the early phase, the subsequent training process exhibits significant fluctuations, with occasional steep drops in performance. We attribute this instability to limited data coverage caused by the near-deterministic nature of the policy and the noise in the environment, which together present a substantial challenge for robust learning.
6.3. OOD generalization of DAMO
Model-based methods are generally more capable of performing OOD generalization than model-free methods. As DAMO follows the idea of behavior regularization, which potentially hinders the agent from exploration, it remains unclear whether DAMO preserves the superior generalization ability Lyu et al. (2022); Liu et al. (2024). To clarify this problem, we evaluate DAMO in the halfcheetah-jump task Yu et al. (2020) where the agent solves a task that entails the effective generalization outside the scope of the offline data. Specifically, a behavior policy is trained online to control the cheetah to move forward as fast as possible. The rewards of the collected trajectories are then reshaped to encourage the cheetah jumping during offline training. Since the behavior policy is trained for a different purpose, the dataset only contains the suboptimal policy for jumping. Therefore, OOD generalization is crucial for the agent to acquire a high-return policy. We compare DAMO with other model-based methods and behavior-regularization methods. The implementation of the experiment and introduction of baselines are presented in Appendix C.5. As shown in Table 4, DAMO achieves comparable performance with model-based methods and outperforms all behavior-regularization methods. This suggests that combining MBRL with behavior regularization enables the agent to achieve both offline conservatism and OOD generalization when handling OOD issues in offline RL.
@@ 6.3. OOD generalization of DAMO
Figure 5. (a) The tuning experiment of $\alpha$ was performed on the hopper-medium-expert dataset for validation. (b) The fixing experiment of $\alpha$ was performed on the halfcheetah-medium-expert dataset for validation.
6.4. Effects of Coefficient $\alpha$
@@
DAMO is regulated by a key hyperparameter: the hyperparameter $\alpha$, which controls the extent of policy discrepancies. While we consider $\alpha$ as a hyperparameter controlling the degree of conservatism, the empirical results in Fig. 5(a) demonstrate that DAMO maintains robust performance across a wide range of $\alpha$ values, with its primary impact observed in the final stage performance. DAMO achieves superior performance with larger $\alpha$ values (e.g., $\alpha=3.0,5.0$), whereas insufficiently small values (e.g., $\alpha=1.0$) tend to promote excessive risk-taking during exploration, failing to converge to an optimal policy. To enhance training stability, we initially follow the prior work Luo et al. (2024a) and fix $\alpha$ to 1.0 in the actor training. We present the concrete implementation in Appendix C.4. However, experimental results demonstrate that DAMO achieves competitive performance even without this specific implementation. To further validate this observation, a comparative experiment is conducted, evaluating both the standard DAMO implementation and its variant without fixed $\alpha$ under their optimal hyperparameter configurations. As shown in Fig. 5(b), both settings achieve competitive performances, confirming that DAMO maintains robust performance regardless of the specific implementation.
7. Conclusion and Discussion
@@
This paper examines limitations in existing offline MBRL approaches, highlighting the crucial need to address behavior inconsistency caused by policy discrepancies between biased models and real environments. Building upon our analysis, we introduce a unified framework that concurrently ensures both synthetic-offline data compatibility and model-environment policy consistency, with experimental validation demonstrating its effectiveness. Significantly, the maximin optimization framework of DAMO demonstrates the benefits of aligning the actor and critic training objectives within the actor-critic architecture, ensuring essential conservatism during policy updates. Our future work could extend DAMO to areas such as offline-to-online RL Li et al. (2023), where the method with adaptability and the capability to bridge the synthetic-to-real distribution gap is in great demand. The limitations of DAMO are focused on two-fold: the sampling of the initial state and the computation of the data alignment term. Although we assume the distribution of the initial state is accessible, some offline datasets do not explicitly label the initial states. Therefore, we randomly sample a state from these offline datasets to serve as the initial state, which creates a deployment gap. A classifier is employed to assist computation. However, the accuracy of the classifier is critical: imprecise classifiers fail to estimate the modified reward, leading to unstable policy training and potentially invalidating the objective as a lower bound.
@@ Appendix A Proofs in the Main Text
We present a detailed proof for each theorem discussed above. The sequence of proofs follows the logic of derivation rather than strictly adhering to their order of appearance in the main text.
A.1. Proof of Theorem 3
To prove Theorem 3, we first introduce an upper bound of the KL-divergence.
Lemma. 0.
For any convex function $f(x)$ satisfying the inequality: $f(x)\geq x\log x,\forall x>0$, the KL-divergence admits the following upper bound:
$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})\leq\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
@@ A.1. Proof of Theorem 3
$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ $=\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]$
$=\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}\cdot\dfrac{\rho_{M}^{\pi}}{\rho_{M}^{\pi}}]$
$=\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
$\leq\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
To establish the final inequality, we invoke the property that the f-divergence dominates the KL divergence, i.e., $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})\geq D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$, which stems from the inequality $f(x)\geq x\log x$ imposed on the convex function $f(x)$. ∎
Deploying the Lemma 4 into the objective $\max\limits_{\pi}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]-\alpha D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$, then we have a lower bound of vanilla RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]$.
Corollary. 0.
We have the following surrogate objective that serves as a lower bound of $\mathcal{J}(\pi)$:
$\mathcal{J}(\pi)\geq\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).\quad\forall\alpha\geq 0$
@@
The following objective is a lower bound of $\mathcal{J}(\pi)$ due to the non-negativity of KL-divergence:
$\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]-\alpha D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$
We substitute the $D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ with its upper bound in theorem 4, then for any policy $\pi$, we have:
We have now completed the proof of Theorem 3.
A.2. Proof of Theorem 1
Next, we present the proof of Theorem 1. We first introduce two lemmas.
Firstly, we define the conjugate functional and provide a corresponding proposition of biconjugate.
Lemma. 0.
Given a functional $f:\mathbb{E}\rightarrow\mathbb{R}$, $\mathbb{E}$ is a Hilbert space. Its Fenchel conjugate functional $f_{\star}:\mathbb{E}\rightarrow\mathbb{R}$ is defined as:
$f_{\star}(y)=\sup\limits_{x\in\mathbb{E}}\{\langle y,x\rangle-f(x)\}$
$\langle y,x\rangle$ is the inner dot of $y,x$. We call the functional $f_{\star\star}=(f_{\star})_{\star}$ as the biconjugate functional of $f$. If $f$ is a convex, closed functional, then its biconjugate functional is itself:
$f_{\star\star}=f$
@@ A.2. Proof of Theorem 1
The proof of the proposition about the biconjugate functional can be found in Fenchel (2014)
Then, in the second lemma, we consider the convex functional.
Lemma. 0.
We can represent f-divergence in a functional form with $x(z)$: $D_{f}(x\|p)=\mathbb{E}_{z\sim p}f(\frac{x(z)}{p(z)})\triangleq h_{p}(x)$, that means we fix the distributional function $p$ and treat $D_{f}(x\|p)$ as a functional $h_{p}(x)$. Then its Fenchel conjugate function is $\mathbb{E}_{z\sim p}f_{\star}(y(z))$
@@
We use $g_{\star}(y)$ to denote the conjugate functional of $D_{f}(x\|p)$, $\mathcal{X}$ to denote the $L^{2}space$. Then we have:
$g_{\star}(y)$ $=\sup\limits_{x\in\mathcal{X}}\{\langle y,x\rangle-h_{p}(x)\}$
$=\sup\limits_{x\in\mathcal{X}}\{\int_{R}y(z)x(z)dz-\int_{R}p(z)\cdot f(\frac{x(z)}{p(z)})dz\}$
$=\sup\limits_{x\in\mathcal{X}}\{\int_{R}p(z)[y(z)\frac{x(z)}{p(z)}-f(\frac{x(z)}{p(z)})]dz\}$
$=\int_{R}p(z)\sup\limits_{x(z)}[y(z)\frac{x(z)}{p(z)}-f(\frac{x(z)}{p(z)})]dz$
$=\int_{R}p(z)\sup\limits_{x^{\prime}(z)}[y(z)x^{\prime}(z)-f(x^{\prime}(z))]dz$
$=\int_{R}p(z)f_{\star}(y(z))dz$
$=\mathbb{E}_{z\sim p}f_{\star}(y(z))$
The first equality is the definition of the Fenchel conjugate function. In the second equality, we give the explicit expressions for the inner product and the KL divergence. In the fifth equality, we make the variable change and substitute $(x(z)/p(z))$ with $x^{\prime}(z)$. The sixth equality follows the definition of the Fenchel conjugate function. ∎
We can employ Lemma 6, 7 to represent the f-divergence in surrogate objective (1) in a MAX-form.
Corollary. 0.
The f-divergence $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ satisfy
$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})=\max\limits_{y(s,a,s^{\prime})}\mathbb{E}_{\rho_{T}^{\pi}}[y(s,a,s^{\prime})]-\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))].$
@@
We can verify that $D_{f}(x\|p)$ is a convex, closed functional. Then the biconjugate functional of $D_{f}(x\|p)$ is itself according to Lemma 6. We keep the notation and use $g_{\star}(y)$ to denote the conjugate function of $D_{f}(x\|p)$, $y(s,a,s^{\prime})$ is any function defined in the transition space. Then $D_{f}(x\|p)$ is the conjugate of $g_{\star}(y)$. Therefore, $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ can be represented by the conjugate functional of $g_{\star}(y)$:
$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ $=\max\limits_{y}\langle y,\rho^{\pi}_{T}\rangle-g_{\star}(y)$
$=\max\limits_{y}\int\rho^{\pi}_{T}\cdot y\ \mathrm{d}s\ \mathrm{d}a\ \mathrm{d}s^{\prime}-g_{\star}(y)$
$=\max\limits_{y}\mathbb{E}_{\rho_{T}^{\pi}}[y]-\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y)]$
The first equality is the definition of the Fenchel conjugate function. In the second equality, we express the inner product in the form of an integral. In the third equality, we represent $g_{\star}(y)$ in the form of expectation according to Lemma 7. ∎
With the Corollary 8, we can attain a maximin form of surrogate objective (1).
Theorem. 9.
Optimizing surrogate objective (1) is equivalent to solving the following maximin problem
$\max\limits_{\pi}\min\limits_{y}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y)]$ (5)
Proof.
Substituting the $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ in surrogate objective (1) with its MAX-form in Corollary 8, then we attain the equivalent maximin form of surrogate objective (1). ∎
Finally, we should make the maximin objective feasible in the offline RL setting, which means that we should eliminate the expectation over $\rho^{\pi}_{T}$ in (5).
Theorem. 10.
The maximin objective (5) is equivalent to the following objective
$\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$ (6)
Proof.
Defining $y(s,a,s^{\prime})=\frac{1}{\alpha}r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\gamma\mathbb{E}_{a^{\prime}\sim\pi}Q(s^{\prime},a^{\prime})-Q(s,a)$. For the convenience, we define $\beta_{t}(s,a,s^{\prime})$
@@
According to definition, $\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}$. Similarly, we can define $\beta_{t}(s,a)$ and $\beta_{t}(s,a,s^{\prime},a^{\prime})$.
$\beta_{t}(s,a)=Pr[s_{t}=s,a_{t}=a|s_{0}\sim\mu_{0},a_{t}\sim\pi,s_{t}\sim T]$
$\beta_{t}(s,a,s^{\prime},a^{\prime})=Pr[s_{t}=s,a_{t}=a,s_{t+1}=s^{\prime},a_{t+1}=a^{\prime}|\mu_{0},\pi,T]$
Substituting $y(s,a,s^{\prime})$ with $r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\gamma\mathbb{E}_{a^{\prime}\sim\pi}Q(s^{\prime},a^{\prime})-Q(s,a)]/\alpha$, then the first term in objective (5) can be represented as:
$\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]$
$=$ $\mathbb{E}_{\rho_{T}^{\pi}}[Q(s,a)-\gamma\mathbb{E}_{a^{\prime}\sim\pi}Q(s^{\prime},a^{\prime})]$
@@
$=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
$=$ $(1-\gamma)[\sum\limits_{s,a}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a)Q-\sum\limits_{s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$